OpenStudy (1dfanatic):

write an equation for the line that is perpendicular to the given line and that passes through the given point. (12,-6); y=4x+1

OpenStudy (anonymous):

do you know relation ship b/w horizontal and perpendicular line?

OpenStudy (1dfanatic):

no

OpenStudy (mathmale):

You are given the line y = 4x - 1. What is the slope of this line? \[m _{1}=?\]

OpenStudy (1dfanatic):

4x

OpenStudy (anonymous):

ok, i tell you that is\[m _{1}*m _{2}=-1\]where m1 is slope of horizontal line, and m2 is slope of perpendicular line.

OpenStudy (mathmale):

To find the slope of a line perpendicular to y = 4x - 1, we take the "negative reciprocal" of \[m _{1}.\]

OpenStudy (mathmale):

Please note that m is a fixed number (constant), whereas x is a variable. The slope of the given line is not 4x; it's the coefficient of x. Therefore, here, \[m _{1}=?\]

OpenStudy (anonymous):

now slope of given line is\[m _{1}=4\]so, the slope of perpendicular line is\[m _{2}=-(1/4)\]

OpenStudy (anonymous):

now use point slope formula\[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

ok

OpenStudy (mathmale):

@newtonson: your comments are helpful. But do not solve the problem for 1Dfanatic; help that person find his or her own solution.

OpenStudy (1dfanatic):

I don't understand I'M sorry...

OpenStudy (anonymous):

ok i tell you in brief.

OpenStudy (mathmale):

Back to basics. You are given the line y=4x-1. Compare this to y=mx+b. What is the apparent value of m, the slope?

OpenStudy (1dfanatic):

mx would be 4x so the intercept would be 1..

OpenStudy (mathmale):

@1Dfanatic : You are correct in that mx = 4x here. What is the value of the slope, m? While you can find the intercept if you wish, it's not needed here.

OpenStudy (1dfanatic):

what do you mean by m?

OpenStudy (mathmale):

As explained earlier, more than once, m represents the slope of the straight line y=mx+b. b represents the y-intercept. It's essential that you understand the meaning of m and b in the equation y=mx+b.

OpenStudy (mathmale):

You are given the equation of a straight line, y=4x-1. Compare this to the general form y=mx+b. What is the value of the slope, m? What is the value of the y-intercept, b?

OpenStudy (anonymous):



OpenStudy (mathmale):

Compare: y = 4x - 1 y = mx +b y=y -1=b m = ??? m is the slope of the given line.